OpenStudy (anonymous):

Solve 2sinx-√ 3=0....... My test is tomorrow and I don't remeber how to do this.. Please guys help me :(!!!

hero (hero):

\[2\sin(x) - \sqrt{3} = 0\]

hero (hero):

Add \(\sqrt{3}\) to both sides to get: \[2\sin(x) = \sqrt{3}\]

OpenStudy (anonymous):

then divide by 2

OpenStudy (anonymous):

see the answer now?

hero (hero):

Divide both sides by 2 to get: \[\sin(x) = \frac{\sqrt{3}}{2}\]

hero (hero):

Take inverse sine of both sides to isolate \(x\)

OpenStudy (anonymous):

any other questions?

OpenStudy (anonymous):

inverse sin (sqrt(3)/2)

OpenStudy (anonymous):

inverse sin (2) is an error because sin is always between 1 and -1

OpenStudy (anonymous):

sqrt ?? so how I solve this problem :(

OpenStudy (anonymous):

what kind of calculator do you have?

OpenStudy (anonymous):

SHARP advanced D.A.L. (EL-520X)

OpenStudy (anonymous):

does it show inverse trigonometry operations as "sin^(-1)" or as "arcsin"

OpenStudy (anonymous):

sin^-1

OpenStudy (anonymous):

but its the same thing.. the question is, how I do it if its an error for the calculator

hero (hero):

Get a TI-Nspire quick. Throw that Sharp calc in the trash.

OpenStudy (anonymous):

sin^(-1) ((sqrt(3))/2)

OpenStudy (anonymous):

what is sqrt :(

OpenStudy (anonymous):

square root sorry

OpenStudy (anonymous):